1,232 is what percent of 220?
a) 689.92%
b) 560%
c) 68.993%
d) 56%
Help please...

Answer by stanbon(75887)

1,232 is what percent of 220?
--------
1232/220 = x/100
x = 5.6*100 = 560
Ans: 560%
===============
Cheers,
Stan H.